# Fork-Merge Decoding: Enhancing Multimodal Understanding in Audio-Visual Large Language Models

### v1

@@ 5 Related Works
Audio-visual large language models. Building upon the success of LLMs, there has been a surge of interest in extending their capabilities to incorporate audio and visual modalities, with text serving as the central modality. ChatBridge zhao2023chatbridge presents a text-centric modality bridging framework trained on limited paired data, whereas models such as PandaGPT su2023pandagpt , ImageBind-LLM han2023imagebind , and OneLLM han2024onellm leverage unified encoders to accommodate various modalities. Other approaches chen2023vast ; lu2024unified ; lyu2023macaw ; panagopoulou2023x ; zhan2024anygpt ; zhang2023video employ modality-specific encoders to better capture distinct feature spaces. To enhance spatial-temporal modeling across modalities, CAT ye2024cat introduces a clue aggregator for cross-modal reasoning, VideoLLaMA2 cheng2024videollama utilizes a spatial-temporal convolutional connector for video synchronization, and video-SALMONN sun2024video proposes a multi-resolution causal Q-Former for audio-visual fusion. Meerkat chowdhury2024meerkat further refines multimodal interactions by aligning audio and visual signals at multiple levels through dedicated interaction modules prior to decoding.
Inference-time reasoning enhancement with LLMs. Recent efforts have explored inference-time strategies to enhance the reasoning capabilities of LLMs without additional training. Chain-of-Thought (CoT) guides LLMs to produce intermediate reasoning steps, and has been extended to VLMs through structured textual representations himakunthala2023let ; ni2024visual ; zhu2023chatgpt or modular reasoning pipelines sun2025mm ; xu2024llava ; yang2023mm . These approaches improve interpretability and robustness through a more explicit reasoning process. In parallel, Contrastive Decoding (CD) improves inference-time decoding by comparing token-level logits between a weaker and a stronger model li2022contrastive . DoLA chuang2023dola develops this idea by contrasting early and late layer outputs within a single model to refine predictions. Recently, VCD leng2024mitigating extends CD to VLMs by injecting Gaussian noise into image inputs and contrasting the resulting biased predictions with the original outputs. Other CD-based approaches kim2024code ; wang2024mitigating enhance decoding robustness by utilizing self-descriptions or distorting instructions. SID huo2024self further advances this line of work by preserving the least informative visual tokens and contrasting their influence on predictions.
@@ 6 Conclusion
We analyze modality bias in current AV-LLMs, where jointly processing audio and visual inputs can hinder balanced reasoning. To address this, we propose Fork-Merge Decoding (FMD)—a simple, training-free, efficient, and model-agnostic inference strategy that separates modality specific understanding in the early decoder layers (the fork phase) and merges their representations in later layers (the merge phase). FMD consistently improves performance on tasks requiring integrated multimodal understanding, as demonstrated across three audio-visual benchmarks using two representative AV-LLMs: VideoLLaMA2 and video-SALMONN. Our approach is broadly applicable to AV-LLMs, offering a plug-and-play solution that enables deeper unimodal and multimodal understanding during inference. We hope this work encourages further research into interpretable reasoning strategies tailored to the unique challenges of MLLMs operating over complex, multi-sensory modalities.
@@
[…]
Appendix A Architectural Analysis of Recent AV-LLMs
We investigate recently proposed AV-LLMs by categorizing them based on open-source availability, the presence of pretrained checkpoints, decoder architecture, and the method used for audio-visual integration (AV fusion). As shown in Table A.1, all models except for the commercial AV-LLMs, GPT-4o [17] and Gemini 1.5 Pro [39], are open-sourced. While a variety of decoder architectures are employed across models, most approaches rely on either token-wise or channel-wise concatenation to fuse audio and visual tokens as input. This suggests that the proposed FMD method has the potential to be applied in a generalizable manner.
Table A.1: Comparison of representative AV-LLMs in terms of open-source availability, checkpoint release, decoder backbone, and fusion strategy. Notably, several models provide full access to code and checkpoints, while others (e.g., GPT-4o and Gemini 1.5 Pro) remain closed. Fusion types are categorized as either token-wise or channel-wise.
Model Open-source Checkpoints Decoder AV fusion
GPT-4o [17] (Unknown) (Unknown)
Gemini 1.5 Pro [39] (Unknown) (Unknown)
Meerkat [7] LLaMA2-7B-Chat Token-wise
video-SALMONN [35] Vicuna-7B-v1.5 Vicuna-13B-v1.5 LLaMA2 Channel-wise
VideoLLaMA-2 [6] Qwne2-7B-Instruct Token-wise
CAT [47] LLaMA2-7B Token-wise
AnyGPT [49] LLaMA2-7B Token-wise
Unified-IO 2 [27] UIO-2-1.1B UIO-2-3.2B UIO-2-6.8B Token-wise
OneLLM [11] LLaMA2-7B Token-wise
X-instructBLIP [32] Vicuna-7B-v1.1 Vicuna-13B-v1.1 Token-wise
ImageBind-LLM [12] LLaMA-7B Channel-wise
Macaw-LLM [28] LLaMA-7B Token-wise
VideoLLaMA [50] Vicuna-7B Vicuna-13B Token-wise
VAST [5] Vicuna-13B Token-wise
PandaGPT [34] Vicuna-13B-v0 Channel-wise
ChatBridge [53] Vicuna-13B Token-wise
Appendix B Further Analysis
@@
Figure A.1: Attention weight analysis in VideoLLaMA2 [6] on the AVHBench dataset [37]. We analyze 100 samples and examine the attention weights in decoder layers after $L_{\text{fork}}$, focusing on the final token. FMD narrows the gap between audio and video attention weights, encouraging more balanced contributions from both modalities.
B.1 Analysis of attention weights beyond the merge point
@@
Through the main paper Figure 1, we demonstrate that in the final layer of VideoLLaMA2 [6], the attention weights are higher for the video input compared to the audio input. To analyze the weight distribution trend across the entire model—not just the final layer—we visualize the average attention weights after the merge point at $L_{\text{fork}}$ in Figure A.1. Specifically, we examine the last token in the sequence and compute the aggregated attention over video and audio segments, averaged across all attention heads, to estimate each modality’s contribution to the prediction. The proposed FMD consistently promotes a more balanced attention distribution between modalities.
(a) Performance on 500 samples from the AVHBench dataset.
(b) Performance on 1000 samples from the AVHBench dataset.
Figure A.2: Layer-wise accuracies of VideoLLaMA2 [6] on each task in the AVHBench [37] dataset. The consistent trends observed across 500 and 1000 samples further validate the 200-sample analysis presented in Figure 5.
Figure A.3: Cosine similarity comparison on VideoLLaMA2 [6] using 100 samples from the AVHBench [37] dataset. We compute the cosine similarity between the final-layer hidden states of the intact input and those of audio- or video-perturbed inputs. The results indicate that video signals are not effectively isolated by additive Gaussian noise, whereas zero-masking reliably suppresses both modalities.
B.2 Layer-wise performances on various dataset sizes
[…]
B.3 Comparison between gaussian noise addition and zero-out masking
@@
We analyze two input perturbation strategies that aim to suppress modality-specific information by replacing the original input: Gaussian noise injection, as proposed in VCD [20], and zero-out masking, as employed in our method. Specifically, we apply each perturbation method to the video and audio inputs, respectively, and compute the cosine similarity between the final-layer hidden states of the perturbed inputs and those of the original inputs.
@@
Appendix C More Qualitative Results
In addition to the qualitative analysis presented in Section 4.3, we include further examples that illustrate both successful outcomes and failure cases, providing a more comprehensive understanding of the model’s behavior.
C.1 Positive cases
@@
C.2 Negative cases
Although our FMD significantly enhances multimodal understanding without requiring additional training, it occasionally produces inaccurate phrases alongside otherwise detailed and accurate descriptions. For example, in the top case of Figure A.8, FMD successfully captures both visual and audio content, whereas vanilla decoding describes only the visual scene. However, the phrase “talking to them” is contextually inappropriate, as “them” refers to the ducks, while the audio indicates the man is actually speaking toward the cameraman. In the bottom example, the mention of “camera flashes” is inaccurate, since only stage lighting is present. While such instances reflect occasional hallucinations, FMD generally produces rich and informative responses by effectively modeling cross-modal relationships. Moreover, at a broader scale, it tends to reduce hallucinations across datasets, thereby improving overall performance, as demonstrated in Table 1. Further investigation into minimizing fine-grained hallucinations remains an important direction for future work.
Appendix D Computational Resource
@@
Appendix E Limitations and Future Works
Our proposed FMD is a simple yet effective inference-time method that requires no additional training. However, its performance inherently depends on the quality and capacity of the underlying model. As a future direction, we aim to develop models that explicitly incorporate unimodal interactions during training. We believe that such an approach will enable the model to more effectively capture and integrate modality-specific information, thereby enhancing its performance on multimodal reasoning tasks.
Appendix F Social Impact
@@
Figure A.4: Qualitative results for audio-driven video hallucination tasks using VideoLLaMA2 [6]. Compared to vanilla decoding, FMD generates more accurate responses by effectively leveraging both audio and visual modalities.
Figure A.5: Qualitative results for video-driven audio hallucination tasks using VideoLLaMA2 [6]. FMD produces more accurate responses in most cases, including instances where visual signals lead to confusion, as shown in the third example.
Figure A.6: Qualitative results for audio-visual matching tasks using VideoLLaMA2 [6]. FMD generates correct responses even when the synthesized audio and video inputs are semantically unrelated.
Figure A.7: Qualitative results for audio-visual description tasks using VideoLLaMA2 [6]. FMD effectively describes both audio and visual content, capturing fine-grained details—such as “A person touches the edge of a wooden clock” in the top example.
Figure A.8: Failure case analysis for audio-visual description tasks using VideoLLaMA2 [6]. While FMD occasionally generates inaccurate details, it still captures both audio and visual content more effectively—and with finer granularity—compared to vanilla decoding.

### v2

@@ 5 Related Works
Audio-visual large language models. Building upon the success of LLMs, there has been a surge of interest in extending their capabilities to incorporate audio and visual modalities, with text serving as the central modality. ChatBridge (Zhao et al., 2023b) presents a text-centric modality bridging framework trained on limited paired data, whereas models such as PandaGPT (Su et al., 2023), ImageBind-LLM (Han et al., 2023), and OneLLM (Han et al., 2024) leverage unified encoders to accommodate various modalities. Other approaches (Chen et al., 2023b; Lu et al., 2024; Lyu et al., 2023; Panagopoulou et al., 2023; Zhan et al., 2024; Zhang et al., 2023) employ modality-specific encoders to better capture distinct feature spaces. To enhance spatial-temporal modeling across modalities, CAT (Ye et al., 2024) introduces a clue aggregator for cross-modal reasoning, VideoLLaMA2 (Cheng et al., 2024) utilizes a spatial-temporal convolutional connector for video synchronization, and video-SALMONN (Sun et al., 2024) proposes a multi-resolution causal Q-Former for audio-visual fusion. Meerkat (Chowdhury et al., 2024) further refines multimodal interactions by aligning audio and visual signals at multiple levels through interaction modules prior to decoding. Recently, Qwen2.5-Omni (Xu et al., 2025) is introduced as a model that perceives audio-visual inputs, generates text and natural speech, and achieves superior performance in audio-visual tasks.
Inference-time reasoning enhancement with LLMs. Recent efforts have explored inference-time strategies to enhance the reasoning capabilities of LLMs without additional training. Chain-of-Thought (CoT) guides LLMs to produce intermediate reasoning steps, and has been extended to VLMs through structured textual representations (Himakunthala et al., 2023; Ni et al., 2024; Zhu et al., 2023a) or modular reasoning pipelines (Sun et al., 2025a; Xu et al., 2024; Yang et al., 2023). These approaches improve interpretability and robustness through a more explicit reasoning process. In parallel, Contrastive Decoding (CD) improves inference-time decoding by comparing token-level logits between a weaker and a stronger model (Li et al., 2023b). DoLA (Chuang et al., 2024) develops this idea by contrasting early and late layer outputs within a single model to refine predictions. Recently, VCD (Leng et al., 2024b) extends CD to VLMs by injecting Gaussian noise into image inputs and contrasting the resulting biased predictions with the original outputs. Other CD-based approaches (Kim et al., 2024; Wang et al., 2024) enhance decoding robustness by utilizing self-descriptions or distorting instructions. SID (Huo et al., 2025) further advances this line of work by preserving the least informative visual tokens and contrasting their influence on predictions.
@@ 6 Conclusion
We analyze modality bias in current AV-LLMs, where jointly processing audio and visual inputs can hinder balanced reasoning. To address this, we propose Fork-Merge Decoding (FMD), a simple, training-free, efficient, and model-agnostic inference strategy that separates modality-specific understanding in the early decoder layers (the fork phase) and merges their representations in later layers (the merge phase). FMD consistently improves performance on tasks requiring integrated multimodal understanding, as demonstrated across three audio-visual benchmarks using three representative AV-LLMs: VideoLLaMA2, video-SALMONN, and Qwen2.5-Omni. Our approach is broadly applicable to AV-LLMs, offering a plug-and-play solution that enables deeper unimodal and multimodal understanding during inference. We hope this work inspires further research addressing the unique challenges of AV-LLMs in complex, multi-sensory settings.
Limitations. While our method demonstrates strong generalization across datasets without requiring ground-truth labels—relying only on 100 random samples to determine $\alpha$—we note that the optimal value of $\alpha$ can vary depending on model-specific characteristics (e.g., $\alpha=0.8$ for VideoLLaMA2 and $\alpha=0.9$ for Qwen2.5-Omni). This suggests that calibrating $\alpha$ is a lightweight yet necessary step when applying FMD to different architectures, and future work could explore automated strategies to further reduce this effort. In addition, although we identify fork layers that generally perform well across tasks and provide extensive analysis, the optimal choice of fork layer may still vary by task, as illustrated in Figure 6. Investigating adaptive or task-specific fork layer selection strategies therefore represents an interesting direction for future research.
@@
Appendix A Further Analysis
A.1 Analysis of attention weights beyond the merge point
@@
Figure A.1: Attention weight analysis in VideoLLaMA2 on the AVHBench dataset. We analyze 100 samples and examine the attention weights in decoder layers after $L_{\text{fork}}$, focusing on the final token. FMD narrows the gap between audio and video attention weights, encouraging more balanced contributions from both modalities.
@@
As shown in Figure 1 of the main paper, the final layer of VideoLLaMA2 (Cheng et al., 2024) assigns higher attention weights to video inputs compared to audio inputs. To examine attention trends across the entire model, we visualize the average attention weights after the merge point at $L_{\text{fork}}$ in Figure A.1. Specifically, we focus on the last token in the sequence and aggregate attention across layers and heads over video and audio segments to estimate each modality’s contribution to the prediction. The proposed FMD consistently produces a more balanced attention distribution between modalities.
A.2 Layer-wise performance across dataset sizes
In the main paper (see Section 4.4 and Figure 6), we analyze task performance across $L_{\text{fork}}$ using 200 samples from the AVHBench dataset. To examine whether this trend holds with more data, we repeat the analysis with 500 and 1,000 samples per task (see Figure A.2a and Figure A.2b). We observe consistent trends across all sample sizes, demonstrating that $L_{\text{fork}}$ can be reliably validated even with a small number of samples.
A.3 Comparison between gaussian noise addition and zero-out masking
@@
We analyze two input perturbation strategies that aim to suppress modality-specific information by replacing the original input: Gaussian noise injection, as proposed in VCD (Leng et al., 2024b), and zero-out masking, as employed in our method. Specifically, we apply each perturbation method to the video and audio inputs, respectively, and compute the cosine similarity between the final-layer hidden states of the perturbed inputs and those of the original inputs.
@@
(a) Performance on 500 samples from the AVHBench dataset.
(b) Performance on 1,000 samples from the AVHBench dataset.
Figure A.2: Layer-wise accuracies of VideoLLaMA2 on each task in the AVHBench dataset. The consistent trends observed across 500 and 1,000 samples further validate the 200-sample analysis presented in Figure 6.
Figure A.3: Cosine similarity comparison on VideoLLaMA2 using 100 samples from the AVHBench dataset. We compute the cosine similarity between the final-layer hidden states of the intact input and those of audio- or video-perturbed inputs. The results indicate that video signals are not effectively isolated by additive Gaussian noise, whereas zero-masking reliably suppresses both modalities.
A.4 Analysis of inference speed across $L_{fork}$
To better understand the relationship between inference speed and the fork layer $L_{fork}$, we evaluate FMD on VideoLLaMA2 using 100 randomly selected AVHBench samples across different fork layer settings.
As shown in Table A.1, deeper fork layers require more decoder layers to be forwarded, leading to longer inference time. At the same time, accuracy drops significantly because later forking disrupts the pretrained model attribute as illustrated in Figure 5 and Figure 6. Based on this, we set $L_{fork}=4$ in VideoLLaMA2 (one out of every seven layers), which yields consistent performance gains while increasing inference time by only 26%.
Table A.1: Analysis between inference speed and fork layer $L_{fork}$ in VideoLLaMA2 on AVHBench. Deeper fork layers require more decoder computation, increasing latency while significantly reducing accuracy as illustrated in Figure A.2. Setting $L_{fork}=4$ (one-seventh of the 28 layers) achieves consistent performance gains with only 26% increase in inference time.
$L_{fork}$ 0 (Vanilla) 1 4 5 10 15 20 25
Latency (sec/token) 0.34 0.40 0.43 0.45 0.51 0.56 0.62 0.68
Relative 1 1.18 1.26 1.32 1.5 1.65 1.82 2
Appendix B More Qualitative Results
In addition to the qualitative analysis presented in Section 4.3, we include further examples that illustrate both successful outcomes and failure cases, providing a more comprehensive understanding of the behavior of the model.
B.1 Positive cases
@@
B.2 Negative cases
Although our FMD significantly enhances multimodal understanding without requiring additional training, it occasionally produces inaccurate phrases alongside otherwise detailed and accurate descriptions. For example, in the top case of Figure A.8, FMD successfully captures both visual and audio content, whereas vanilla decoding describes only the visual scene. However, the phrase “while she does it” is temporally incorrect, since the speaker talks before blow-drying. In the bottom example, the mention of “cup” is inaccurate, as no cup is present. Although such cases reflect occasional misunderstandings, FMD generally produces richer and more informative responses by modeling cross-modal relationships (e.g., capturing “talking”). Moreover, at a broader scale, it tends to reduce hallucinations across datasets, thereby improving overall performance, as demonstrated in Table 1. Further investigation into minimizing fine-grained hallucinations remains an important direction for future work.
Appendix C Computational Resource
@@
Appendix D The Use of LLMs
We use LLMs to refine words and sentences for a formal academic writing style and to identify relevant related work. For evaluation, we employ LLMs to assess long-form responses, such as AV captioning from AVHBench, following the protocol of VideoLLaMA2.
Appendix E Social Impact
@@
Figure A.4: Qualitative results for audio-driven video hallucination tasks using VideoLLaMA2. Compared to vanilla decoding, FMD generates more accurate responses by effectively leveraging both audio and visual modalities.
Figure A.5: Qualitative results for video-driven audio hallucination tasks using VideoLLaMA2. FMD produces more accurate responses in most cases, including instances where visual signals lead to confusion, as shown in the third example.
Figure A.6: Qualitative results for audio-visual matching tasks using VideoLLaMA2. FMD generates correct responses even when the synthesized audio and video inputs are semantically unrelated.
Figure A.7: Qualitative results for audio-visual description tasks using VideoLLaMA2. FMD effectively describes both audio and visual content, capturing fine-grained details (bold)—such as “A man is talking about the bus” in the top example.
Figure A.8: Failure case analysis on audio-visual description tasks with VideoLLaMA2 (Cheng et al., 2024). FMD may produce occasional errors (red), but captures both audio and visual content more effectively and with finer-grained details (bold) than vanilla decoding.